OpenStudy (anonymous):

Aaron run 6 miles in 48 minutes, Dario ran 5 miles in 35 minutes and Serge ran 7 miles in 56 minutes. Who ran the fastest per mile? please help I'm kind of confused

OpenStudy (anonymous):

\[\{\text{Aaron},\text{Darlo},\text{Serge}\}\to \left\{\frac{6}{48},\frac{5}{35},\frac{7}{56}\right\}\to \left\{\frac{1}{8},\frac{1}{7},\frac{1}{8}\right\} \]Darlo because he can run a mile in 7 minutes. The other two runners require 8 minutes. Do you have any questions regarding the data organization process?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

and no